OpenStudy (anonymous):

The region R is bounded by the x-axis, x = 1, x = 3, and y = 1/x^3.

OpenStudy (anonymous):

A. Find the area R. B. Find the value of h, such that the vertical line x = h divides the region R into two Regions of equal area.

OpenStudy (agent0smith):

Oh no i think you prob just made a mistake in calculating. \[\Large \frac{-1 }{ 2(3)^2 } - \frac{ -1 }{ 2(1)^2 } = \]

OpenStudy (anonymous):

Ohhh, I see where I went wrong. I only used x = 3 and forgot to use x = 1. What about number 2? I'm not particular sure how to approach that problem.

OpenStudy (agent0smith):

\[\Large \int\limits_{1}^{h}\frac{ 1 }{x^3 } dx = \int\limits_{h}^{3}\frac{ 1 }{x^3 } dx\]integrate and solve for h.

OpenStudy (anonymous):

Oh sorry, didn't realize you were back! Would the answer be 0.22?

OpenStudy (agent0smith):

Idk, too tired to check, i might tomorrow.

OpenStudy (anonymous):

Okay, no problem. :D